OpenStudy (anonymous):

Find a point-slope form for the equation of the line satisfying the conditions. Slope -4, passing through (3,9)

OpenStudy (anonymous):

How would I approach this I understand how the equation should look like in slope-intercept form.

OpenStudy (anonymous):

plug in to y-y1=m(x-x1)

OpenStudy (anonymous):

y-9=-4(x-3)

OpenStudy (anonymous):

You got it